OpenStudy (anonymous):

Thom worked 18 hours last week at the sporting goods store and earned $92.00. If he continues to earn the same hourly pay, how many additional hours must he work to earn another $138.00?

OpenStudy (anonymous):

can someone please try to help im not good with proportions

OpenStudy (anonymous):

Well, if he worked 18 hours, and got payed 92 dollars, how much did he get payed per hour?

OpenStudy (anonymous):

The question is how much did he make per hour.... per means to divide so word for word make per hour means 92 / 18 That would give you 5.11 per hour. so how many $5.11 are there in $138 138/5.11 27.0058 hours I would assume you would answer 27 hours.

OpenStudy (anonymous):

ok yea that makes sense so just break it down step by step

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

how do i find a proportion for x like this x/6 = 6/4

OpenStudy (anonymous):

but its like x is above the 6 and the 6 is above the 4 i guess like cross multiply

OpenStudy (anonymous):

Whenever you have a fraction = a fraction you can always cross multiply (top left times bottom right = bottom left times top right) x 6 - = - 6 4 so x times 4 = 6 times 6 4x = 36 x = 9

OpenStudy (anonymous):

wow thank u i should of typed it like that lol ok so 4x = 36 and x = 9 but is that the answer ?

OpenStudy (anonymous):

Yes.. you are wanting to know what x equals and x = 9

OpenStudy (anonymous):

ok so i should only get one answer which is x

OpenStudy (anonymous):

In other words 9 6 - = - 6 4 if you reduce both of those fractions you get 3 - 2

OpenStudy (anonymous):

ok can u please tell me how to reduce do u just find the factors that go into each numbers

OpenStudy (anonymous):

ok im confused im trying to do 9.5/19 and x/30

OpenStudy (anonymous):

please can someone help

OpenStudy (anonymous):

i did 9.5 times 30 and got 285 but now im stuck

OpenStudy (anonymous):

9.5 x --- = --- 19 30 30(9.5) = 19x 285 = 19x x = 15

OpenStudy (anonymous):

oh so i do 30 times 9.5 ok sorry i did the oposite

OpenStudy (anonymous):

It doesn't matter which way you multiply, you get the same answer.

OpenStudy (anonymous):

oh ok yea because i did get 285

OpenStudy (anonymous):

At a certain time of day, a tree casts a 9-foot shadow at the same time that a person 6 feet tall casts a 4-foot shadow. What is the height of the tree?

OpenStudy (anonymous):

shadow / height = shadow / height

OpenStudy (anonymous):

9 divided by 6 = 1.5 correct

OpenStudy (anonymous):

9 6 - = - x 4 9(4) = 6x 36 = 6x 6 = x

OpenStudy (anonymous):

i had to cross mutiply

OpenStudy (anonymous):

WAIT I WAS WRONG.

OpenStudy (anonymous):

um ok

OpenStudy (anonymous):

9 4 - = - x 6 9(6) = 4x 54 = 4x 16 = x

OpenStudy (anonymous):

Shadow of tree Shadow of person ------------ = -------------- Height of tree Height of person

OpenStudy (anonymous):

ok i understand now so that would mean the answer is 16=x

OpenStudy (anonymous):

OR You can do Shadow of tree Height of tree ------------- = ------------ Shadow of person Height of person 9 x - = - 4 6 9(6) = 4x 54 = 4x x = 13.5

OpenStudy (anonymous):

I am SO sorry... don't know what is wrong with my math. They both = 13.5 because 54/4 = 13.5 (I don't know WHERE I got 16)

OpenStudy (anonymous):

its ok really everyone makes masteakes

OpenStudy (anonymous):

ok so the answer is 13.5

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

:) thank you wonderful

OpenStudy (anonymous):

U R Welcome

OpenStudy (anonymous):

im going to try the next one if i get stuck ill get back to to

OpenStudy (anonymous):

x/27 = 4/9 i did x times 9 and i got 9x the i did 9x times 9x and got 81 is that correct

OpenStudy (anonymous):

blexting

OpenStudy (anonymous):

am i correct blexting